# Near-Optimal Sample Complexities of Divergence-based S-rectangular Distributionally Robust Reinforcement Learning

### v1

@@ C.1 Proof of Proposition 2
By definition and $|\sup_{X}f-\sup_{X}g|\leq\sup_{X}|f-g|$, we have
$P\left(\left|\hat{\mathbf{T}}^{*}(v)(s)-\mathcal{T}^{*}(v)(s)\right|>t\right)$
$=P\left(\left|\sup_{\pi\in\Pi}\hat{\mathbf{T}}^{\pi}(v)(s)-\sup_{\pi\in\Pi}\mathcal{T}^{\pi}(v)(s)\right|>t\right)$
$\leq P\left(\sup_{\pi\in\Pi}\left|\hat{\mathbf{T}}^{\pi}(v)(s)-\mathcal{T}^{\pi}(v)(s)\right|>t\right).$
@@
$\left|\hat{\mathbf{T}}^{\pi}(v)(s)-\mathcal{T}^{\pi}(v)(s)\right|$ $\leq\left|\sup_{\lambda>0}f(P_{s,n},R(s,\cdot,\cdot)+\gamma v,\lambda)-\sup_{\lambda>0}f(P_{s},R(s,\cdot,\cdot)+v,\lambda)\right|$ (C.2)
$\leq\sup_{\lambda>0}\left|f(P_{s},R(s,\cdot,\cdot)+\gamma v,\lambda)-f(P_{s},R(s,\cdot,\cdot)+v,\lambda)\right|.$
@@
$P\left(\left|\hat{\mathbf{T}}^{*}(v)(s)-\mathcal{T}^{*}(v)(s)\right|>t\right)\leq P\left(\sup_{\lambda>0,d_{a}\in\Delta(\mathcal{A})}\left|f(P_{s},R(s,\cdot,\cdot)+\gamma v,\lambda)-f(P_{s},R(s,\cdot,\cdot)+\gamma v,\lambda)\right|\right)$
@@
$P\left(\left\|\hat{\mathbf{T}}^{*}(v)-\mathcal{T}^{*}(v)\right\|_{\infty}>t\right)$
$\leq P\left(\sup_{s}\sup_{\lambda\geq 0,d\in\Delta(\mathcal{A})}|f(P_{s,n},R(s,\cdot,\cdot)+\gamma v,\lambda)-f(P_{s},R(s,\cdot,\cdot)+v,\lambda)|>t\right)$
@@
$P\left(\left\|\hat{\mathbf{T}}^{*}(v)-\mathcal{T}^{*}(v)\right\|_{\infty}>t\right)\leq\eta.$
@@ C.2 Proof of Lemma 6
$=\left|\sum_{a\in\mathcal{A}}\lambda\frac{m_{a,n}[e^{d_{a}(\|u\|_{L^{\infty}(\mu_{a})}-u)/\lambda}-1]}{\mu_{a,n}(t)[e^{d_{a}(\|u\|_{L^{\infty}(\mu_{a})}-u)/\lambda}]}\right|.$
@@
[…]
@@ C.4 Proof of Theorem 1
with probability $1-\eta$. ∎
@@ D.1 Proof of Proposition 3
$P\left(\left|\hat{\mathbf{T}}^{*}(v)(s)-\mathcal{T}^{*}(v)(s)\right|>t\right)$
$\leq$ $P\left(\sup_{\pi}\left|\hat{\mathbf{T}}^{\pi}(v)(s)-\mathcal{T}^{\pi}(v)(s)\right|>t\right)$
$\leq$ $P\left(\sup_{d\in\Delta(|\mathcal{A}|)}\gamma\left|\sup_{\boldsymbol{\eta}\in\mathbb{R}^{|\mathcal{A}|}}f(\boldsymbol{\mu}_{n},R(s,\cdot,\cdot)+\gamma v,\boldsymbol{\eta})-\sup_{\boldsymbol{\eta}\in\mathbb{R}^{|\mathcal{A}|}}f(\boldsymbol{\mu},R(s,\cdot,\cdot)+\gamma v,\boldsymbol{\eta})\right|>t\right).$
@@
$P\left(\sup_{d\in\Delta(|\mathcal{A}|)}\gamma\left|\sup_{\boldsymbol{\eta}\in\mathbb{R}^{|\mathcal{A}|}}f(\boldsymbol{\mu}_{n},u,\boldsymbol{\eta})-\sup_{\boldsymbol{\eta}\in\mathbb{R}^{|\mathcal{A}|}}f(\boldsymbol{\mu},u,\boldsymbol{\eta})\right|>t\right)$
@@
$P\left(\left\|\hat{\mathbf{T}}^{*}(v)-\mathcal{T}^{*}(v)\right\|_{\infty}>t\right)$
$\leq P\left(\sup_{s\in\mathcal{S}}\gamma\sup_{d\in\Delta(\mathcal{A})}\left|\sup_{\boldsymbol{\eta}\in\mathbb{R}^{|\mathcal{A}|}}f(P_{s,n},R(s,\cdot,\cdot)+\gamma v,\boldsymbol{\eta})-\sup_{\boldsymbol{\eta}\in\mathbb{R}^{|\mathcal{A}|}}f(P_{s},R(s,\cdot,\cdot)+\gamma v,\boldsymbol{\mu})\right|>t\right)$
$\leq 2|\mathcal{S}|^{2}|\mathcal{A}|\exp\left(-\frac{n\mathfrak{p}_{\wedge}}{12}\right)+2|\mathcal{S}|^{2}|\mathcal{A}|\exp\left(-\frac{t^{2}}{2\gamma^{2}}\left(\frac{4c^{2}\|R(s,\cdot,\cdot)+\gamma v\|_{\infty}^{2}}{n\mathfrak{p}_{\wedge}}+\frac{2c\|R(s,\cdot,\cdot)+\gamma v\|_{\infty}t}{3\gamma n\mathfrak{p}_{\wedge}}\right)^{-1}\right).$
@@
$P\left(\left\|\hat{\mathbf{T}}^{*}(v)-\mathcal{T}^{*}(v)\right\|_{\infty}>t\right)\leq\eta,$
@@ D.4 Proof of Theorem 2
with probability $1-\eta$. ∎

### v2

@@ C.1 Proof of Proposition 2
$\sup_{\pi\in\Pi}\left|\hat{\mathbf{T}}^{\pi}(v)(s)-\mathcal{T}^{\pi}(v)(s)\right|$ $=\sup_{d\in\Delta(\mathcal{A})}\left|\sup_{\lambda>0}f(P_{s,n},R(s,\cdot,\cdot)+\gamma v,\lambda)-\sup_{\lambda>0}f(P_{s},R(s,\cdot,\cdot)+\gamma v,\lambda)\right|$ (C.2)
$\leq\sup_{\lambda>0,d\in\Delta(\mathcal{A})}\left|f(P_{s,n},R(s,\cdot,\cdot)+\gamma v,\lambda)-f(P_{s},R(s,\cdot,\cdot)+\gamma v,\lambda)\right|.$
@@
$P\left(\sup_{\pi\in\Pi}\left|\hat{\mathbf{T}}^{\pi}(v)(s)-\mathcal{T}^{\pi}(v)(s)\right|>t\right)\leq P\left(\sup_{\lambda>0,d_{a}\in\Delta(\mathcal{A})}\left|f(P_{s},R(s,\cdot,\cdot)+\gamma v,\lambda)-f(P_{s},R(s,\cdot,\cdot)+\gamma v,\lambda)\right|\right)$
@@
$P\left(\sup_{\pi\in\Pi}\left\|\hat{\mathbf{T}}^{\pi}(v)-\mathcal{T}^{\pi}(v)\right\|_{\infty}>t\right)$
$\leq P\left(\sup_{s}\sup_{\lambda\geq 0,d\in\Delta(\mathcal{A})}|f(P_{s,n},R(s,\cdot,\cdot)+\gamma v,\lambda)-f(P_{s},R(s,\cdot,\cdot)+\gamma v,\lambda)|>t\right)$
@@
$P\left(\sup_{\pi\in\Pi}\left\|\hat{\mathbf{T}}^{\pi}(v)-\mathcal{T}^{\pi}(v)\right\|_{\infty}>t\right)\leq\eta.$
Since $|\sup_{X}f-\sup_{X}g|\leq\sup_{X}|f-g|$,
$P\left(\sup_{\pi\in\Pi}\left\|\hat{\mathbf{T}}^{*}(v)-\mathcal{T}^{*}(v)\right\|_{\infty}>t\right)$ $=P\left(\left\|\sup_{\pi\in\Pi}\hat{\mathbf{T}}^{\pi}(v)-\sup_{\pi\in\Pi}\mathcal{T}^{\pi}(v)\right\|_{\infty}>t\right)$
$\leq P\left(\sup_{\pi\in\Pi}\left\|\hat{\mathbf{T}}^{\pi}(v)-\mathcal{T}^{\pi}(v)\right\|_{\infty}>t\right)$
$\leq\eta.$
@@ C.2 Proof of Lemma 8
$\left|\sum_{a\in\mathcal{A}}\lambda\frac{m_{a,n}[e^{d_{a}(\|u\|_{L^{\infty}(\mu_{a})}-u)/\lambda}-1]}{\mu_{a,n}(t)[e^{d_{a}(\|u\|_{L^{\infty}(\mu_{a})}-u)/\lambda}]}\right|.$
@@
Applying this result and $\left|\sum\cdot\right|\leq\sum|\cdot|$, we obtain
@@ C.4 Proof of Theorem 1
with probability $1-\eta$.
The preceding step provides a uniform bound on the value-function estimation error. We now turn to the policy suboptimality gap. We first introduce following lemma.
Lemma 10.
Let $\hat{\pi}^{*}:=\operatorname*{argmax}_{\pi}V_{\hat{\mathcal{P}}}^{\pi}$, and we have
$0\leq\sup_{\pi\in\Pi}V_{\mathcal{P}}^{\pi}-V_{\mathcal{P}}^{\hat{\pi}^{*}}\leq 2\sup_{\pi\in\Pi}\left\|V_{\hat{\mathcal{P}}}^{\pi}-V_{\mathcal{P}}^{\pi}\right\|_{\infty}$
According to Corollary 1, Lemma 10 and Proposition 2, we have
$\sup_{\pi\in\Pi}V_{\mathcal{P}}^{\pi}-V_{\mathcal{P}}^{\hat{\pi}^{*}}$ $\leq 2\sup_{\pi\in\Pi}\left\|V_{\hat{\mathcal{P}}}^{\pi}-V_{\mathcal{P}}^{\pi}\right\|_{\infty}$
$\leq\frac{2}{1-\gamma}\sup_{\pi\in\Pi}\left\|\hat{\mathbf{T}}^{\pi}(V_{\mathcal{P}}^{\pi})-\mathcal{T}^{\pi}(V_{\mathcal{P}}^{\pi})\right\|_{\infty}$
$\leq\frac{18\|R+\gamma v\|_{\infty}}{\sqrt{n\mathfrak{p}_{\wedge}}(1-\gamma)}\sqrt{\log{(4|\mathcal{S}|^{2}|\mathcal{A}|/\eta)}}$
$\leq\frac{18}{\sqrt{n\mathfrak{p}_{\wedge}}(1-\gamma)^{2}}\sqrt{\log{(4|\mathcal{S}|^{2}|\mathcal{A}|/\eta)}}$
C.5 Proof of Lemma 10
$\sup_{\pi\in\Pi}V_{\mathcal{P}}^{\pi}-V_{\mathcal{P}}^{\hat{\pi}^{*}}$ $=\sup_{\pi\in\Pi}V_{\mathcal{P}}^{\pi}-\sup_{\pi\in\Pi}V_{\hat{\mathcal{P}}}^{\pi}+V_{\hat{\mathcal{P}}}^{\hat{\pi}^{*}}-V_{\mathcal{P}}^{\hat{\pi}^{*}}$
$\leq\left\|\sup_{\pi\in\Pi}V_{\mathcal{P}}^{\pi}-\sup_{\pi\in\Pi}V_{\hat{\mathcal{P}}}^{\pi}\right\|_{\infty}+\left\|V_{\hat{\mathcal{P}}}^{\hat{\pi}^{*}}-V_{\mathcal{P}}^{\hat{\pi}^{*}}\right\|_{\infty}$
$\leq 2\sup_{\pi\in\Pi}\left\|V_{\mathcal{P}}^{\pi}-V_{\hat{\mathcal{P}}}^{\pi}\right\|_{\infty}$
@@ D.1 Proof of Proposition 3
$P\left(\sup_{\pi}\left|\hat{\mathbf{T}}^{\pi}(v)(s)-\mathcal{T}^{\pi}(v)(s)\right|>t\right)$
$\leq P\left(\sup_{d\in\Delta(|\mathcal{A}|)}\left|\sup_{\boldsymbol{\eta}\in\mathbb{R}^{|\mathcal{A}|}}f(\boldsymbol{\mu}_{n},R(s,\cdot,\cdot)+\gamma v,\boldsymbol{\eta})-\sup_{\boldsymbol{\eta}\in\mathbb{R}^{|\mathcal{A}|}}f(\boldsymbol{\mu},R(s,\cdot,\cdot)+\gamma v,\boldsymbol{\eta})\right|>t\right).$
@@
$P\left(\sup_{d\in\Delta(|\mathcal{A}|)}\left|\sup_{\boldsymbol{\eta}\in\mathbb{R}^{|\mathcal{A}|}}f(\boldsymbol{\mu}_{n},u,\boldsymbol{\eta})-\sup_{\boldsymbol{\eta}\in\mathbb{R}^{|\mathcal{A}|}}f(\boldsymbol{\mu},u,\boldsymbol{\eta})\right|>t\right)$
@@
$P\left(\sup_{\pi\in\Pi}\left\|\hat{\mathbf{T}}^{\pi}(v)-\mathcal{T}^{\pi}(v)\right\|_{\infty}>t\right)$
$\leq P\left(\sup_{s\in\mathcal{S}}\sup_{d\in\Delta(\mathcal{A})}\left|\sup_{\boldsymbol{\eta}\in\mathbb{R}^{|\mathcal{A}|}}f(P_{s,n},R(s,\cdot,\cdot)+\gamma v,\boldsymbol{\eta})-\sup_{\boldsymbol{\eta}\in\mathbb{R}^{|\mathcal{A}|}}f(P_{s},R(s,\cdot,\cdot)+\gamma v,\boldsymbol{\eta})\right|>t\right)$
$\leq 2|\mathcal{S}|^{2}|\mathcal{A}|\exp\left(-\frac{n\mathfrak{p}_{\wedge}}{12}\right)+2|\mathcal{S}|^{2}|\mathcal{A}|\exp\left(-\frac{t^{2}}{2}\left(\frac{4c^{2}\|R(s,\cdot,\cdot)+\gamma v\|_{\infty}^{2}}{n\mathfrak{p}_{\wedge}}+\frac{2c\|R(s,\cdot,\cdot)+\gamma v\|_{\infty}t}{3\gamma n\mathfrak{p}_{\wedge}}\right)^{-1}\right).$
@@
$P\left(\sup_{\pi\in\Pi}\left\|\hat{\mathbf{T}}^{\pi}(v)-\mathcal{T}^{\pi}(v)\right\|_{\infty}>t\right)\leq\eta,$
Since $|\sup_{X}f-\sup_{X}g|\leq\sup_{X}|f-g|$,
$P\left(\sup_{\pi\in\Pi}\left\|\hat{\mathbf{T}}^{*}(v)-\mathcal{T}^{*}(v)\right\|_{\infty}>t\right)$ $=P\left(\left\|\sup_{\pi\in\Pi}\hat{\mathbf{T}}^{\pi}(v)-\sup_{\pi\in\Pi}\mathcal{T}^{\pi}(v)\right\|_{\infty}>t\right)$
$\leq P\left(\sup_{\pi\in\Pi}\left\|\hat{\mathbf{T}}^{\pi}(v)-\mathcal{T}^{\pi}(v)\right\|_{\infty}>t\right)$
$\leq\eta.$
@@ D.4 Proof of Theorem 2
with probability $1-\eta$. The preceding step provides a uniform bound on the value-function estimation error. We now turn to the policy suboptimality gap. According to Corollary 2, Lemma 10 and Proposition 3, we have
$\sup_{\pi\in\Pi}V_{\mathcal{P}}^{\pi}-V_{\mathcal{P}}^{\hat{\pi}^{*}}$ $\leq 2\sup_{\pi\in\Pi}\left\|V_{\hat{\mathcal{P}}}^{\pi}-V_{\mathcal{P}}^{\pi}\right\|_{\infty}$
$\leq\frac{2}{1-\gamma}\sup_{\pi\in\Pi}\left\|\hat{\mathbf{T}}^{\pi}(V_{\mathcal{P}}^{\pi})-\mathcal{T}^{\pi}(V_{\mathcal{P}}^{\pi})\right\|_{\infty}$
$\leq\frac{6\cdot 2^{k^{*}}k^{*}\|R+\gamma v\|_{\infty}}{\sqrt{n\mathfrak{p}_{\wedge}}(1-\gamma)}\sqrt{\log\left(4|S|^{2}|\mathcal{A}|/\eta\right)}$
$\leq\frac{6\cdot 2^{k^{*}}k^{*}}{\sqrt{n\mathfrak{p}_{\wedge}}(1-\gamma)^{2}}\sqrt{\log\left(4|S|^{2}|\mathcal{A}|/\eta\right)}$